OpenStudy (anonymous):

The theoretical probability of undesirable side effects resulting from taking Grebex is 2 in 19. If 380 people take Grebex to lower their blood pressure, how many will encounter undesirable side effects?

OpenStudy (anonymous):

2/19 * 380 = 40

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

im not sure how you do these types of problems.

OpenStudy (anonymous):

2 in 19 in this context means that if you take 19 people then we would expect 2 people to get side effects. So we calculate 2/19 ths of 380. with fractions 'of' means multiply. 2/19 times 380 = (2 * 380) / 19 = 760 / 19 = 40

OpenStudy (anonymous):

ok....

OpenStudy (anonymous):

sure?

OpenStudy (anonymous):

yea.. so if i have a circle and they want a probability of the chance there is 2 yellow slices what would the probability be it landing on yello

OpenStudy (anonymous):

its split into 8 slices. and 2 slices are yellow

OpenStudy (anonymous):

it would depend on the size of the slices if for example there were 2 yellow slices and three red all the same size and they filled the circle then the chance of landing on yellow = no. of yellow slices / total no. of slices = 2/5

OpenStudy (anonymous):

oh ok then in that case, the probability = 2 /10 or 1/5

OpenStudy (anonymous):

so it would be 2/8 chances of landing on the yellow

OpenStudy (anonymous):

sorry yes 2/8 not 2/10 as total no. slices = 8

OpenStudy (anonymous):

u got that right.

OpenStudy (anonymous):

yea im not as stupid as i feel... lol

OpenStudy (anonymous):

lol